OpenStudy (anonymous):

Simplify. –24 – 8 ÷ (–4) A. –8 B. –22 C. –26 D. 8

OpenStudy (anonymous):

@ObeyFCS

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

@YoungStudier

OpenStudy (anonymous):

Using PEMDAS –24 – 8 ÷ (–4)=-22

OpenStudy (anonymous):

wtf i have it

OpenStudy (anonymous):

you didnt explain, so no

OpenStudy (anonymous):

-8/-4=2 24-2=-22

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@ObeyFCS @ObeyFCS

OpenStudy (anonymous):

C

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

-8/-4 =2 then -24 -2 = -26

OpenStudy (anonymous):

WOOLAH xD